E11-19 (Depletion Computations—Timber) Stanislaw Timber Company owns 9,000 acres of timberland
purchased in 2003 at a cost of $1,400 per acre. At the time of purchase, the land without the timber was
valued at $400 per acre. In 2004, Stanislaw built fire lanes and roads, with a life of 30 years, at a cost of
$84,000. Every year, Stanislaw sprays to prevent disease at a cost of $3,000 per year and spends $7,000 to
maintain the fire lanes and roads. During 2005, Stanislaw selectively logged and sold 700,000 board feet of
timber, of the estimated 3,500,000 board feet. In 2006, Stanislaw planted new seedlings to replace the trees
cut at a cost of $100,000.
Instructions
(a) Determine the depreciation expense and the cost of timber sold related to depletion for 2005.
(b) Stanislaw has not logged since 2005. If Stanislaw logged and sold 900,000 board feet of timber in
2016, when the timber cruise (appraiser) estimated 5,000,000 board feet, determine the cost of timber
sold related to depletion for 2016.
Please explain the answers so I can fully understand the process thank you
Answer #1

a.

Cost of Timberland            1,400.00
Cost of Land                400.00
Cost of Timber            1,000.00
Cost of 9,000 Timber    9,000,000.00
Estimated board fee    3,500,000.00
Total Cost per board ft                       2.57
Board feet Timber Sold          700,000.00
Cost of timber sold to depletion    1,800,000.00
Fire lanes and road Life 30 Yrs
Fire lanes and road Cost          84,000.00
Depreciation expense            2,800.00

b. Cost of timber sold = 9,000,000 - 2,100,000 +100,000 = 7,000,000

Depletion = 7,000,000/5,000,000 x 900,000 = 1,260,000